Living matter has an activity of 260 Bq/kg due to carbon – 14. If a sample of wood from a
burial site has an activity of 155 Bq/kg, estimate the age of the site. (Assume the half – life
of carbon – 14 is 5730 years).
1
Expert's answer
2020-06-30T18:26:42-0400

 The activity of the wood decreases as:


"A_t=A_0(\frac{1}{2})^\frac{t}{h}""T_\frac{1}{2}=5730"

"A_0=260bq\/kg"

A=155dq-Activity at the time T

Therefore;

"A=0_0(\u2212\\frac{T}{T}\\frac{\u200b1}{2}\u2192T=T_\\frac{1}{2}log_2[A_\\frac{0}{A}]"

"5730\times log_2[\frac{260}{155}=4276years"

Therefore the age of the site is 4276 years.